# LLMs for Time Series: an Application for Single Stocks and Statistical Arbitrage

### v1

@@ Abstract
Recently, LLMs (Large Language Models) have been adapted for time series prediction with significant success in pattern recognition. However, the common belief is that these models are not suitable for predicting financial market returns, which are known to be almost random. We aim to challenge this misconception through a counterexample. Specifically, we utilized the Chronos model from Ansari et al. (2024) and tested both pretrained configurations and fine-tuned supervised forecasts on the largest American single stocks using data from Guijarro-Ordonnez et al. (2022). We constructed a long/short portfolio, and the performance simulation indicates that LLMs can in reality handle time series that are nearly indistinguishable from noise, demonstrating an ability to identify inefficiencies amidst randomness and generate alpha. Finally, we compared these results with those of specialized models and smaller deep learning models, highlighting significant room for improvement in LLM performance to further enhance their predictive capabilities.
@@ 1 Introduction
LLMs (Large Language Models) gained widespread popularity when ChatGPT convinced many that machines were intelligent enough to reason like humans, even though the underlying techniques merely determine the most likely sequences of words that could respond to a prompt. The introduction of the transformer architecture by Vaswani et al. (2017) was a key development, as it enabled fast training on large datasets. LLMs are composed of many layers of transformers and have around a billion parameters. The T5 (Text-To-Text Transfer Transformer) architecture was introduced by Raffel et al. (2023), and further advanced the field from T5-small size with 60 million of parameters to T6-11B with 11 billion of parameters). Amatrianin (2023) describes the catalog of LLMs models from Albert to ChatGPT and classifies T5 among others.
@@
Garza et al. (2023) introduce TimeGPT, the first foundation model for time series, capable of generating accurate predictions for diverse datasets not seen during training. They evaluate their pre-trained model against established statistical, machine learning, and deep learning methods, demonstrating that TimeGPT zero-shot inference excels in performance. Ansari et al. (2024) also sought to adapt these highly efficient LLMs to time series forecasting. Their approach involved representing real numbers in different bins (using a vocabulary of 4096 tokens) and training a T5 architecture on a wide range of time series data (around 90 billion observations). They produced several pretrained models of varying sizes, ranging from tiny (11 millions of parameters) to large. TimesFM (Time Series Foundation Model) is another pretrained time-series foundation model developed by Das et al. (2024). Rasul et al. (2024) uses a lagged features for tokenization. Nie et al. (2023); Ekambaram et al. (2023) are also applying LLM to multi time series. Motivated by recent advances in large language models for Natural Language Processing, Ansari et al. (2024); Das et al. (2024) designed a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Their models are based on pretraining a patched-decoder style attention model on a large time-series corpus, and can work well across different forecasting history lengths, prediction lengths and temporal granularities.
Brugiere and Turinici (2024) tested transformers for financial time series and showed that the algorithm cannot predict returns, but can only predict squared returns. Konigstein (2024) stated that LLMs present challenges, but also opportunities, particularly for long-term financial time series forecasting.
@@
Deep learning has become very popular among researchers in finance, both for asset pricing and systematic strategies: Guijarro-Ordonnez et al. (2022) implemented transformers (with only 769 parameters) but coupled them with convolutional layers and tested them on quantitative trading strategies applied to single stocks. Their results, without accounting for trading costs, were encouraging. Guijarro-Ordonnez et al. (2022) applied deep learning algorithms to residual daily returns after removing common factors using techniques such as Principal Component Analysis (PCA), Fama-French factors, or Instrumental Principal Component Analysis (IPCA), implementing the methodology from Kelly et al. (2019), who used financial data to constrain the eigenvectors. This approach was also developed and justified by Valeyre (2019). Transformers have also been used by Jiang et al. (2023) to identify patterns in images for time series forecasting. Wood et al. (2022) applied deep learning techniques with only a few parameters to identify the most effective trend-following indicators enhancing and timing trend-following strategies. Transformers have also been used to extract common returns, as demonstrated by Gu et al. (2021). Qyrana (2024) applied a simple STR factor to residual returns using an autoencoder-based factor model, subsequently generating a highly profitable trading strategy. Chen et al. (2021) used a deep learning technique to identify anomalies in asset pricing,
@@
Our goal, in this study, is to evaluate deep learning algorithms with more than 11 million parameters for forecasting financial returns. In contrast, studies such as Chen et al. (2021), Jiang et al. (2023), Wood et al. (2022), Guijarro-Ordonnez et al. (2022), and Brugiere and Turinici (2024) utilized models with at most a few hundred parameters so their models were not really ”deep learning” models. This limitation was due to their inability to pre-train models with millions of parameters using large datasets from outside the financial industry.
@@
So we first conducted a zero-shot evaluation of the predictions from pretrained and fine-tuned supervised time series foundation LLMs Chronos by Ansari et al. (2024), which were pretrained on 13 datasets that do not include single stock data or stock indices, using the datasets of the residual returns of American single stocks published by Guijarro-Ordonnez et al. (2022). The interest in using only zero-shot evaluation is that it provides more convincing results, as overfitting is less likely in this case. Indeed overfitting is a major concern in machine learning when applied to trading, which often makes honest results appear suspect. Another interesting aspect is demonstrating how the algorithm can adapt and display ’intelligence’ in trading without being specifically trained for that purpose. We aim to simulate a portfolio that goes long on positive predictions and short on negative predictions.
Secondly, we seek to compare these results with well-known standard Short Term Reversal (STR) or trend-following approaches, as described in Jegadeesh (1990); Jegadeesh and Titman (1993).
@@ 2.1 Chronos as the LLM model
The model used was ”amazon/chronos-t5-tiny” version of the chronos with 11 million of parameters which was pretrained by Ansari et al. (2024) on 14 datasets (Brazilian cities temperatures, Mexico city bikes, Solar, Spanish energy and weather, Taxi, USHCN, weatherbench, wiki daily, wind farms) but not on financial time series. The model has 11 millions of parameters.
We used a ’context’ period of 100 days so that Chronos guess the next day, knowing only the previous 100 days. We focused only on the next day return forecast. We used 100 days as a compromise to avoid running out of memory while giving Chronos a chance to capture some patterns. Additionally, we decided to limit the study to predicting the next daily returns. Although we could have considered predicting the next weekly or monthly returns, we believed it would be easier for Chronos to capture patterns over a very short-term horizon.
@@
We adjusted the 11 million weights of the ”amazon/chronos-t5-tiny” model through training (fine-tuning) using our datasets, setting $\tau$, the maximum training steps, to 5, 15, or 40. Training was conducted daily during the backtest, using the data available on each respective date and starting from the weights of the previous day.
The details of the parameters for both the pretrained case and the fine tuning are described in the appendix B.1 and B.2.
2.2 Financial market time series as data
We used 3 different datasets of the residual daily returns released by Guijarro-Ordonnez et al. (2022), derived from their analysis of securities in the CRSP dataset from January 1978 to end 2016. Their focus was on the most liquid stocks to mitigate trading and friction issues. Specifically, they considered stocks whose market capitalization in the prior month exceeded 0.01% of the total market capitalization of that month, resulting in a selection of approximately the largest 550 stocks on average. Guijarro-Ordonnez et al. (2022) released their datasets of residual returns on GitHub. We utilized their three datasets, each corresponding to their standard parameters with $K=5$ (i.e. with 5 factors):
@@ 2.2 Financial market time series as data
  • IPCA factors with a rolling windows of 240 months with the details described in Kelly et al. (2019)
  • PCA factors with a rolling windows of 252 days
  • FF factors (Fama-French 3 factor model+ investment and profitability factors) with a rolling windows of 60 days
@@
$K=5$ appears to be the optimal according to Guijarro-Ordonnez et al. (2022) when applying their convolution and transformer process with a gross sharpe ratio of 3.21 for Fama-French, 3.36 for the PCA and 4.16 for the IPCA. Nevertheless the sharpe in net appears to be significant only before 2006.
Using residual returns allows for a less correlated dataset, which is crucial for deep learning. We can note that Gu et al. (2021) for example also used the same three different datasets to test their model.
@@ 2.3 Description of the different simulated strategies
Our experiment was organized in three parts.
2.3.1 Strategy based on the forecast of the zero-shot version of Chronos
@@
  • Computing $\hat{\chi}_{d,i}$ , the Chronos average prediction of the next daily return, derived in the Eq 2 conditioned to a rolling window of the last 100 days. In Eq 2 we used the average of different ”equiweight” scenarii $\chi$ computed by Chronos of $\hat{r}_{d+1}$ knowing only $r_{d,i}...r_{d-99,i}$. We used two possible inputs for Chronos:
    • Either the last 100 residual daily returns $r_{d,i}...r_{d-99,i}$ of the single stock $i$, when $\alpha=0$ in Eq 1.
    • or the last $\hat{r}_{d,i}...\hat{r}_{d-99,i}$ the exponential moving average of the last 100 residual daily returns derived in Eq 1 with $\alpha>0$. We tested $\alpha$ values of 0.1, 0.2, 0.3, 0.4, 0.5, and 0.8, all different from zero. This option allows the model to account for the well-known weak negative autocorrelation of daily returns, potentially improving its forecasting ability. However, in this case, the model must outperform the Short-Term Reversal strategy described in Eq. 10.
  • Predicting of the next returns with $\tilde{\chi}_{d,i}$ which is derived in Eq 3.
  • Calculating the weights of the portfolio $\hat{\omega}$ derived in Eq 6 which ranks through $\Re=\text{ArgSort}$ every day $d$ the different $\tilde{\chi}_{d,i}$. In this method the median rank is withdrawn through $\frac{N}{2}$ where $N$ is the number of stocks. A normalization of the weights is derived in Eq 6 to target a gross investment of 1. This process ensures that the portfolio is 50% long and 50% short every day, with weights proportional to the distance in ranking from the median stock according to $\tilde{\chi}$. Valeyre (2019) proved that this approach is the mathematically optimal method and better than just buying the top quintile and short the bottom quintile. A ’resized’ version is also tested when the weights are also inversely proportional to the volatility as derived in Eq 5 where $\sigma$ are the standard deviation of the daily returns on the previous 100 days and ${\mathbf{M}}$ is the median.
  • Simulating, ${\mathcal{P}}_{d+1}$, the performance of the portfolio for the next day through Eq 8. We then reconstructed the cumulative returns and calculated the gross Sharpe ratio, excluding any trading costs. We also simulate $\left[{\mathcal{P}}_{d+1}\right]$ for the resized version through Eq 9.
@@
${\hat{\omega}^{\chi}}_{d,i}=\frac{{\omega^{\chi}}_{d,i}}{\sum_{i}{\left|{\omega^{\chi}}_{d,i}\right|}}$ (6)
$\left[{\hat{\omega}^{\chi}}_{d,i}\right]^{r}=\frac{\left[{\omega^{\chi}}_{d,i}\right]^{r}}{\sum_{i}{\left|\left[{\omega^{\chi}}_{d,i}\right]^{r}\right|}}$ (7)
@@
2.3.2 Strategy based on the forecast of the fine tuned version of Chronos

### v2

# Large Language Models for Time Series: an Application for Single Stocks and Statistical Arbitrage

@@ Abstract
Large Language Models (LLMs) have been adapted for time series prediction with significant success in pattern recognition. However, the common belief is that these models are not suitable for predicting financial market returns, which are known to be almost random. We aim to challenge this misconception through a counterexample. Specifically, we utilize the Chronos model from Ansari et al. (2024) and test both pretrained configurations and fine-tuned supervised forecasts on the largest American single stocks using data from Guijarro-Ordonnez et al. (2021). We construct a long/short portfolio, and the performance simulation indicates that LLMs can in reality handle time series that are nearly indistinguishable from noise, demonstrating an ability to identify inefficiencies amidst randomness and generate alpha. Finally, we compare these results with benchmark models, highlighting significant room for improvement in LLM performance to further enhance their predictive capabilities.
@@ 1 Introduction
Large Language Models (LLMs) gained widespread popularity when ChatGPT convinced many that machines were intelligent enough to reason like humans, even though the underlying techniques merely determine the most likely sequences of words that could respond to a prompt. Besides, with regards to the integration of generative AI in the financial industry, Citadel CEO Ken Griffin says ”the tech hasn’t done much for hedge funds where it matters most— beating the market” [^1] https://blinks.bloomberg.com/news/stories/T48R8PBQ99TS. This study challenges this view by applying for the first time a LLM to a large portfolio of U.S stocks. Recently, the introduction of the transformer architecture by Vaswani et al. (2017) was a key development, as it enabled fast training on large datasets. LLMs are composed of many layers of transformers and have around a billion parameters. For instance, the T5 (Text-To-Text Transfer Transformer) architecture was introduced by Raffel et al. (2023), and further advanced the field from T5-small size with 60 million of parameters to T6-11B with 11 billion of parameters). Amatrianin (2023) describes the catalog of LLMs models from Albert to ChatGPT and classifies T5 among others.
The incorporation of Transformers into trading, risk, and the portfolio management industry is on the verge of transforming the financial landscape due to the oncoming technological gap. The reason is straightforward: They are reaching the state-of-the-art performance of learning from time series in many domains including Traffic, Climate, Energy, and recently, Finance. The transformer is indeed a deep learning architecture that enhances considerably the analysis and processing of huge volume of diverse and complex data that are valuable for businesses and decision-makers on a real time basis. For instance, it helps making timely decision of buying and selling securities or of optimizing portfolio with risk prediction. LLMs, such like Generative Pre-Trained Transformers, are designed for specific tasks useful in the financial industry like real time translation or sentiment analysis. Still, the most challenging, crucial and valuable task remains the time series forecasting. Time series in finance are difficult to predict because they are modeled as random walk process based on a sequence of random observations generated by a stochastic process over time. In addition, financial time series are subject to noise, temporal dependence, non-stationarity and sometimes missing values. In an efficient market it should be nearly impossible to do prediction. In the real world, there exist thousands of financial time series since several decades, but most of them are not liquid enough to be included in a sample for prediction. For example, small capitalization shares or penny stocks could be subject to less frequent or stale quotes. Econometrics models (e.g. ARIMA) are well suited for predicting financial future outcomes because they provide a clear representation of the system, which is very useful in an industry where regulators require explanations. Unfortunately, they suffer from limitations when it comes to handling large and complex data structure. Significant advancements were made from the machine learning models (e.g. Nearest-Neighbor methods, Shrinkage methods, Tree-based methods and Boosting, Support Vector Machines classifiers, Neural Networks approaches) that proved successful in capturing complex data set. On top of that, deep learning models (e.g. Long Short-Term Memory) went even further through modelling heterogeneous and long-term data to extract patterns from complex data structure, which enhances forecasting accuracy. However, they are exposed to a sort of performance saturation because they are relatively small models with hundreds of parameters. Transformers address this limitation allowing for billions of parameters to be estimated on large scale dataset (please see Wen et al. 2024 for a survey on transformers on time series). They are reshaping the field of time series forecasting given their empirical success in the so-called “one-model-for-all-datasets” (Wu et al. (2025)). The vanilla Transformer is built on the encoder-decoder structure where the model architecture is based on an attention mechanism to draw global dependencies between inputs and outputs (see Vaswani et al. (2017)). This encoder-decoder architecture has become a promising tool for predicting complex multivariate time series. LLMs share the same architecture with transformers. LLMs for time series (e.g. BERT) are able to perform time series forecasting across a variety of datasets and are capable of incorporating missing data (Gruver et al. (2023)). The application of such LLMs for time series prediction is an emerging field that excels in predicting time series with clear patterns (Tang et al. (2025)).
@@
Garza et al. (2023) introduce TimeGPT, the first foundation model for time series, capable of generating accurate predictions for diverse datasets not seen during training. They evaluate their pre-trained model against established statistical, machine learning, and deep learning methods, demonstrating that TimeGPT zero-shot inference excels in performance. Ansari et al. (2024) also sought to adapt these highly efficient LLMs to time series forecasting. Their approach involved representing real numbers in different bins (using a vocabulary of 4096 tokens) and training a T5 architecture on a wide range of time series data (around 90 billion observations). They produced several pretrained models of varying sizes, ranging from tiny (11 millions of parameters) to large. TimesFM (Time Series Foundation Model) is another pretrained time-series foundation model developed by Das et al. (2024). Rasul et al. (2024) uses a lagged features for tokenization. Nie et al. (2023) build a transformer model for multivariate time series forecasting by using segmentation of time series into subseries-level patches that served as input tokens to the Transformer and also by using channel-independence for which each channel contains a unique time series that shares the same Transformer weights through all the series. Ekambaram et al. (2023) develop a neural architecture exclusively composed of multi-layer perceptron modules for multivariate forecasting and representation learning on patched time series. This novel MLP-Mixer is also patching-based and it exploits various time-series features for multivariate forecasting. Motivated by recent advances in Large Language Models for Natural Language Processing, Ansari et al. (2024); Das et al. (2024) designed a time-series foundation model for forecasting whose out-of-the-box zero-shot performance on a variety of public datasets comes close to the accuracy of state-of-the-art supervised forecasting models for each individual dataset. Their models are based on pretraining a patched-decoder style attention model on a large time-series corpus, and can work well across different forecasting history lengths, prediction lengths and temporal granularities.
@@
Deep learning has become very popular among researchers in finance, both for asset pricing and systematic strategies: Brugiere and Turinici (2024) tested transformers for financial time series and showed that the algorithm cannot predict returns, but can only predict squared returns. Konigstein (2024) stated that LLMs present challenges, but also opportunities, particularly for long-term financial time series forecasting. Guijarro-Ordonnez et al. (2021) implemented transformers (with only 769 parameters) but coupled them with convolutional layers and tested them on quantitative trading strategies applied to single stocks. Their results, without accounting for trading costs, were encouraging. They applied deep learning algorithms to residual daily returns after removing common factors using techniques such as Principal Component Analysis (PCA), Fama-French factors, or Instrumental Principal Component Analysis (IPCA), implementing the methodology from Kelly et al. (2019), who used financial data to constrain the eigenvectors. This approach was also developed and justified by Valeyre (2019). Transformers have also been used by Jiang et al. (2023) to identify patterns in images for financial price time series. Wood et al. (2022) applied deep learning techniques with only a few parameters to identify the most effective trend-following indicators enhancing and timing trend-following strategies. Transformers have also been used to extract common returns, as demonstrated by Gu et al. (2021). Qyrana (2024) applied a simple Short Term Reversal factor fed by residual returns using an autoencoder-based factor model, subsequently generating a highly profitable trading strategy (i.e. through buying losing stocks based on residual returns and shorting winner stocks). Chen et al. (2021) used a deep learning technique to identify anomalies in asset pricing.
@@
However the question remains whether a Large Language Model optimized for time series outside the field of finance is sufficiently intelligent and sensitive to capture inefficiencies and market anomalies. Despite the burgeoning literature, there are still no papers that apply a deep learning technique with millions of parameters to financial forecasting while this is the natural number of parameters to consider in the context of LLMs. Studies such as Chen et al. (2021), Jiang et al. (2023), Wood et al. (2022), Guijarro-Ordonnez et al. (2021), and Brugiere and Turinici (2024) utilized models with at most a few hundred parameters so their models were not real deep learning models. This limitation was due to their inability to pre-train models with millions of parameters using large datasets from outside the financial industry. This study fills the gap by implementing deep learning algorithms with more than 11 million parameters for forecasting financial returns. This methodological approach attempts to solve the problem to what extent LLMs are intelligent enough to predict financial returns. Therefore, the contribution of this paper is to be the first that use LLMs for financial time series forecasting.
@@
The employed methodology is two-fold. First, we conduct a zero-shot evaluation of the predictions from pretrained and fine-tuned supervised time series foundation LLMs Chronos by Ansari et al. (2024), which were pretrained on 13 datasets that do not include single stock data or stock indices, using the datasets of the residual returns of American single stocks published by Guijarro-Ordonnez et al. (2021). The interest in using only zero-shot evaluation is that it provides more convincing results, as overfitting is less likely in this case. Indeed overfitting is a major concern in machine learning when applied to trading, which often makes honest results appear suspect. Another interesting aspect is demonstrating how the algorithm can adapt and display ’intelligence’ in trading without being specifically trained for that purpose. We simulate a portfolio that goes long the stocks whose predicted next return is positive and short the stocks whose predicted next return is negative. Secondly, we compare these results with the well-known standard Short Term Reversal (STR) or trend-following approaches, as described in Jegadeesh (1990); Jegadeesh and Titman (1993). The main finding reveals that LLMs optimized on non-financial time series are still capable of producing significant forecasts for financial time series.
@@ 2.1 Chronos as the LLM model
Our objective is to assess whether one LLM-based time series model could capture inefficiencies or market anomalies in financial markets.
We limited our test to the Chronos model because it stands out as a best-in-class LLM for time series forecasting, delivering robust performance, particularly in zero-shot and long-horizon forecasting tasks. Its adaptability and scalability make it a valuable tool across a wide range of forecasting applications. The model used was ”amazon/chronos-t5-tiny” version of the chronos which was pretrained by Ansari et al. (2024) on 14 datasets (Brazilian cities temperatures, Mexico city bikes, Solar, Spanish energy and weather, Taxi, USHCN, weatherbench, wiki daily, wind farms) but not on financial time series. The model has 11 millions of parameters. Chronos was presented by Konigstein (2024). Her thorough comparative examination includes several LLM models, such as Lag-Llama from Meta (Rasul et al. (2024)), PatchTST from IBM (Nie et al. (2023)), and TSMixer from IBM (Ekambaram et al. (2023)). According to Konigstein (2024), Chronos was both the easiest to implement and the most effective among the models tested, based on internal evaluations across a variety of models: local models, task-specific models, in-domain pretrained models, and out-of-domain pretrained models. We do not claim that it is optimal or superior to simpler models without pretraining, especially given that out-of-sample results in quantitative finance papers sometimes lack of reliability and sometimes are not confirmed once the paper is published as mentioned by McLean et al. (2016) and Harvey et al. (2016). That is why we limited our investigation to two well established and very simple benchmarks such as both the Short Term Reversal (STR) and ARIMA, while including the machine learning model of Guijarro-Ordonnez et al. (2021) as we used their data. We used a ’context’ period of 100 days so that Chronos guess the next day, knowing only the previous 100 days. We focused only on the next day return forecast. We used 100 days as a compromise to avoid running out of memory while giving Chronos a chance to capture some patterns. Additionally, we decided to limit the study to predicting the next daily returns. While forecasting weekly or monthly returns was a possible alternative, we considered that Chronos would be more effective at identifying patterns over a shorter time horizon.
@@
We used both the zero-shot pretrained and fine-tuned versions of Chronos. To implement the fine-tuned version, we adjusted the 11 million weights of the ’amazon/chronos-t5-tiny’ model by training it on our dataset, setting $\tau$—the maximum number of training steps and one of Chronos’s key parameters—to 5, 15, or 40. Training was conducted daily during the backtest, using the data available on each respective date and starting from the weights of the previous day. The details of the parameters used for both the pretrained and fine-tuned cases are provided in Sections 2.3.1 and 2.3.2, with additional information and the corresponding Python code available in Appendices B.1 and B.2.
2.2 Dataset description
We use the datasets from Guijarro-Ordonnez et al. (2021), who restricted their analysis to the most liquid stocks to mitigate trading and friction issues. Specifically, they considered stocks whose market capitalization in the prior month exceeded 0.01% of the total market capitalization of that month, resulting in a selection of approximately the largest 550 stocks on average. This portfolio of the most liquid U.S. stocks over such a long period appears as the most appropriate sample for a deep learning application. The total number of stocks exceeds 550 because the composition changes daily and we have a total of 9483 different stocks in the whole period. It is very common in financial studies to test strategies on the largest 500 U.S. stocks. Indeed evaluating investment strategies or identifying market anomalies often involves testing on large-cap U.S. stocks, particularly those in the S&P 500 index. This practice is prevalent due to the S&P 500’s representation of a substantial portion of the U.S. equity market and the availability of high-quality data: Hou et al. (2020) examines 452 documented anomalies in finance. The authors construct a vast data library and apply rigorous replication tests, emphasizing the use of NYSE breakpoints and value-weighted returns to mitigate the influence of microcap stocks. Their methodology underscores the importance of focusing on larger, more liquid stocks, such as those in the S&P 500, to ensure robustness in empirical findings. It is worth noting that Gu et al. (2021), in their effort to gather sufficient data to train their 500,000222The figure of 660,000 parameters for their simple version is our own estimation, based on the code released at https://github.com/dkyol/Asset-Pricing-Model. However, this implementation does not appear to be officially validated by Gu et al. (2021) and may not be fully appropriate, as it assigns different weights to each stock, whereas one would typically expect a shared parameter structure across stocks-parameter models, explicitly stated that, ”unlike the existing literature”, they did not exclude micro-cap stocks. They applied no filters, resulting in an average of over 6,200 stocks per day in their dataset. Also as we mainly use the pretrained case of Chronos, we did not need a larger dataset for the zeros shot version or the fine tuned one.
The three different datasets of daily residual returns (IPCA, PCA and FF) provided by Guijarro-Ordonnez et al. (2021) are all based on the same raw universe of securities from the CRSP dataset spanning January 1978 to the end of 2016, but obtained using different methods for extracting the residuals. Guijarro-Ordonnez et al. (2021) released their datasets of residual returns on GitHub. We utilized their three datasets whose statistics are displayed in Tab. 1, each corresponding to their standard parameters with the number of factors, $K$, equal to five:
@@ 2.2 Dataset description
  • IPCA factors with a rolling windows of 240 months with the details described in Kelly et al. (2019). We report the following summary statistics for daily residual returns: a mean of $4.35\times 10^{-6}$ and a standard deviation of $0.0066$.
  • PCA factors with a rolling windows of 252 days. We report the following summary statistics for daily residual returns: a mean of $2.31\times 10^{-6}$ and a standard deviation of $0.0059$.
  • FF factors (Fama-French market, value, size, investment and profitability factors) with a rolling windows of 60 days. We report the following summary statistics for daily residual returns: a mean of $2.95\times 10^{-6}$ and a standard deviation of $0.0068$.
Dataset Mean SD Skewness Kurtosis Total number of residual returns
IPCA $4.35\times 10^{-6}$ $0.0066$ $1.38$ $583$ $4079040$
PCA $2.31\times 10^{-6}$ $0.0059$ $1.16$ $594$ $4079040$
$F$ $2.95\times 10^{-6}$ $0.0068$ $1.16$ $504$ $4079040$
Table 1: Statistics of the daily residual returns of single stocks included in the 3 datasets provided by Guijarro-Ordonnez et al. (2021)
@@
$K=5$ appears to be the optimal hyper parameter according to Guijarro-Ordonnez et al. (2021) when applying their convolution and transformer process with a performance measure, the gross sharpe ratio, of 3.21 for Fama-French, 3.36 for the PCA and 4.16 for the IPCA. Nevertheless the sharpe ratio in net appears to be significant only before 2006. Using residual returns allows for a less correlated dataset, which is crucial for deep learning. It is worth noting that Gu et al. (2021) also used IPCA, PCA and FF as benchmarks to compare the performance of the residuals produced by their autoencoder asset pricing model.
@@ 2.3 Description of the different simulated strategies
Our experiment is organized in three parts. We began by testing the zero-shot version of Chronos, then we implement the fine-tuned version, and finally compared the results against three different benchmarks. We used exactly the same evaluation procedure, described in Section 2.3.1, which constructs the portfolio based on the forecasts generated by each method. Only Guijarro-Ordonnez et al. (2021), which in one of the benchmarks, directly determined the portfolio without going through a separate forecasting step.
2.3.1 Zero-shot version
@@
  • Computing $\hat{\chi}_{d,i}$ , the Chronos average prediction of the next residual daily return, derived in the Eq 2 conditioned to a rolling window of the last 100 days. In Eq 2 we used the average of different ”equiweight” scenarii $\chi$ computed by Chronos of the forecasted residual return, $\hat{r}_{d+1}$, knowing only $r_{d,i}...r_{d-99,i}$. We used two possible inputs for Chronos:
    • Either the last 100 residual daily returns $r_{d,i}...r_{d-99,i}$ of the single stock $i$.
    • or the last $\hat{r}_{d,i}...\hat{r}_{d-99,i}$ the exponential moving average of the last 100 residual daily returns derived in Eq 1 with $\alpha>0$. In fact, the second possibility is merely a more general case of the first, which corresponds to the particular case where $\alpha=0$ in Eq 1. We tested $\alpha$ values of 0.1, 0.2, 0.3, 0.4, 0.5, and 0.8, all different from zero. This option allows the model to account for the well-known weak negative autocorrelation of daily returns, potentially improving its forecasting ability. However, in this case, the model must outperform the Short-Term Reversal strategy described in Eq. 10 as Eq. 1 is close to Eq. 10.
  • Predicting of the next residual returns with $\tilde{\chi}_{d,i}$ which is derived in Eq 3.
  • Calculating the weights of the portfolio $\hat{\omega}$ derived in Eq 6 which uses $\omega^{\chi}_{d}$ derived in Eq 4 which ranks through $\Re=\text{ArgSort}$ every day $d$ the different the forecasted residual return $\tilde{\chi}_{d,i}$ from Eq 3. In this method, in Eq 4, the median rank is withdrawn through $\frac{N}{2}$ where $N$ is the number of stocks. A normalization of the weights is derived in Eq 6 to target a gross investment of 1. This process ensures that the portfolio is 50% long and 50% short every day, with weights proportional to the distance in ranking from the median stock according to the forcasted residual returns $\tilde{\chi}$ derived in Eq 3. Valeyre (2019) proved that this approach is the mathematically optimal method and better than just buying the top quintile and short the bottom quintile as it is invested on a larger number of stocks with a better diversification. A ’resized’ version to manage better risk through reducing weights on volatile stocks is also tested when the weights are also inversely proportional to the volatility as derived in Eq 5 where $\sigma$ are the standard deviation of the daily returns on the previous 100 days and $\mathbf{M}$} is the median.
  • Simulating the performance of the portfolio determined by the weights in Eq 6: We derived ${\mathcal{P}}_{d+1}$, the performance of the portfolio for the next day through Eq 8. We then reconstructed the cumulative returns and calculated the gross Sharpe ratio, excluding any trading costs. We also simulate $\left[{\mathcal{P}}_{d+1}\right]$ for the resized version through Eq 9 which used weights derived in Eq 5.
@@
${\hat{\omega}^{\chi}}_{d,i}=\frac{{\omega^{\chi}}_{d,i}}{\sum_{j}{\left|{\omega^{\chi}}_{d,j}\right|}}$ (6)
$\left[{\hat{\omega}^{\chi}}_{d,i}\right]^{r}=\frac{\left[{\omega^{\chi}}_{d,i}\right]^{r}}{\sum_{j}{\left|\left[{\omega^{\chi}}_{d,j}\right]^{r}\right|}}$ (7)
@@
2.3.2 Fine-tuned version